Leave answers in terms of Pi..
1. Find the circumference of a circle with radius 8?
2. Find the area of a circle with radius 4?

Answer by CharlesG2(828)

Leave answers in terms of Pi..
1. Find the circumference of a circle with radius 8?
C = 2 * R * pi = 2 * 8 * pi = 16pi
2. Find the area of a circle with radius 4?
A = R^2 * pi = 4^2 * pi = 16pi